# Dynamics of the sine-Gordon equation on tadpole graphs

### v1

@@ 2 Well-posedness for the sine-Gordon on a tadpole graph
The proof of Theorem 2.1 requires some preliminary material. We start by describing the spectrum of the one parameter family of self-adjoint operators $(-\Delta,D_{Z})$ for $Z\in\mathbb{R}$ defined in (1.2) and (1.7) (for a related result in the context of tadpole graphs see Noja and Pelinovsky [42], Proposition A.1). In order to simplify the notation we will write $\mathcal{F}_{Z}$ to denote the operator $-\Delta$ acting on $L^{2}(\mathcal{G})$ with domain $D_{Z}$, for each $Z\in\mathbb{R}$.
@@
For every $Z\in\mathbb{R}$ the essential spectrum of the self-adjoint operator $\mathcal{F}_{Z}$ is purely absolutely continuous and $\sigma_{\mathrm{\tiny{ess}}}(\mathcal{F}_{Z})=\sigma_{\mathrm{\tiny{ac}}}(\mathcal{F}_{Z})=[0,\infty)$. If $Z>0$ then $\mathcal{F}_{Z}$ has exactly one negative eigenvalue, i.e., its point spectrum is $\sigma_{\mathrm{\tiny{pt}}}(\mathcal{F}_{Z})=\{-\varrho_{Z}^{2}\}$, where $\varrho_{Z}>0$ is the only positive root of the trascendental equation
@@
$\boldsymbol{\Phi}_{Z}=\begin{pmatrix}\cosh(\varrho_{Z}x/c_{1})\\ e^{-\varrho_{Z}x/c_{2}}\end{pmatrix}.$
@@
First, it is to be observed that from Theorem A.4 and Proposition A.5 the Schrödinger (symmetric) operator $(-\Delta,D(-\Delta))$ with domain $D(-\Delta)$ (which is defined in (A.5)) and acting on $L^{2}(\mathcal{G})$ has deficiency indices $n_{\pm}(-\Delta)=1$ and consequently the one parameter family of self-adjoint extensions $\mathcal{F}_{Z}$ satisfies $n(\mathcal{F}_{Z})\leqq 1$ for all $Z\in\mathbb{R}$.
@@
[…]
@@
assuming $Z>0$. Clearly, $G(0)=-Z<0$ and $G(\varrho)\to\infty$ as $\varrho\to\infty$. Moreover,
@@
for all $\varrho\in(0,\infty)$ and, thus, $G$ is strictly increasing. By the intermediate value theorem there exists a unique value $\varrho_{Z}>0$ such that $G(\varrho_{Z})=0$.
@@
Let $Z\in\mathbb{R}$. For any $\lambda\in\mathbb{C}$ such that $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ we have that $\lambda$ belongs to the resolvent set of $\mathcal{A}=JE$, with $D(\mathcal{A})=D_{Z}\times L^{2}(\mathcal{G})$; moreover, the resolvent,
@@
The proof is very similar to that of Theorem 2.2 in [14], so we gloss over some of the details. Suppose that for $\mathbf{U}=(\boldsymbol{u},\boldsymbol{v})\in H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})$ there holds the relation $(\lambda I_{4}-\mathcal{A})^{-1}=(\boldsymbol{\phi},\boldsymbol{\psi})\in D(\mathcal{A})$. This relation holds if and only if
$\boldsymbol{u}$ $=\lambda\boldsymbol{\phi}-\boldsymbol{\psi},$
$\boldsymbol{v}$ $=\mathcal{F}_{Z}\boldsymbol{\phi}+\lambda\boldsymbol{\psi}.$
Upon substitution we obtain $(\mathcal{F}_{Z}+\lambda^{2}I_{2})\boldsymbol{\phi}=\boldsymbol{v}+\lambda\boldsymbol{u}$. But, by hypothesis, $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ so that $\boldsymbol{\phi}=-R(-\lambda^{2}:\mathcal{F}_{Z})(\boldsymbol{v}+\lambda\boldsymbol{u})$ and, consequently, $\boldsymbol{\psi}=-\lambda R(-\lambda^{2}:\mathcal{F}_{Z})(\boldsymbol{v}+\lambda\boldsymbol{u})-\boldsymbol{u}$, as claimed. ∎
It is actually possible to provide some explicit formulae for $R(-\lambda^{2}:\mathcal{F}_{Z})$ in the case where $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ and $\lambda\in\mathbb{R}$.
@@
Fix $Z\in\mathbb{R}$ and consider $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ with $\lambda\in\mathbb{R}$. For any $\boldsymbol{u}=(u_{1},u_{2})\in L^{2}(\mathcal{G})$ we have:
@@
    In the case when $Z\leq 0$ and for all $\lambda>0$ (without loss of generality) such that $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$, set $\Phi:=(\mathcal{F}_{Z}+\lambda^{2}I_{2})^{-1}\boldsymbol{u}$ to obtain
[…]
@@
    In the case when $Z>0$ and for all $\lambda>0$ (without loss of generality) such that $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$ and with $-\lambda\neq\varrho_{Z}^{2}$, set $\Psi:=(\mathcal{F}_{Z}+\lambda^{2}I_{2})^{-1}\boldsymbol{u}$ to obtain
@@
Formulae (2.6) and (2.7) result from the variation of constants formula: a solution to the homogeneous problem plus a convolution with the Green’s function associated to the operator (see formula (2.6) when $Z\leqq 0$). In the case when there exists a negative eigenvalue (that is, when $Z>0$) one also needs to project onto the associated eigenspace (formula (2.7)). The reader can directly verify that $(\mathcal{F}_{Z}+\lambda^{2}I_{2})\Phi=\boldsymbol{u}$ if $Z\leqq 0$, and that $(\mathcal{F}_{Z}+\lambda^{2}I_{2})\Psi=\boldsymbol{u}$ if $Z>0$, whenever $-\lambda^{2}\in\rho(\mathcal{F}_{Z})$. We have specialized both formulae for $\lambda>0$ for convenience, without loss of generality. These formulae are particular cases of the well-known Krein’s formula for the resolvents of self-adjoint extensions of a symmetric operator (cf. Albeverio et al. [4], Appendix A; Albeverio and Kurasov [5], §1.2). For the Laplacian case see, for example, Theorem 3.1.2 in Albeverio et al. [4], p. 77, or Angulo and Ponce [17], eq. (3.13), for the periodic case. ∎
@@
$\beta:=\begin{cases}\varrho_{Z}^{2},&\text{if }Z>0,\\ 0,&\text{if }Z\leqq 0.\end{cases}$ (2.9)
We will denote by $H^{1}_{Z}(\mathcal{G})$ the space $H^{1}(\mathcal{G})$ with the norm $\|\cdot\|_{X^{1}_{Z}}$. Moreover, the following well-defined inner product in $H^{1}_{Z}(\mathcal{G})$,
[…]
@@
The following theorem verifies that the operator $\mathcal{A}=JE$ is indeed the infinitesimal generator of a $C_{0}$-semigroup upon application of the classical Lumer-Phillips theory.
@@
For convenience of the reader we sketch the main steps of the proof (without lost of generality we consider $c_{j}^{2}=1$):
  1. (1)
    We consider the Hilbert space $X_{Z}\equiv H^{1}_{Z}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})$ with inner product $\langle\cdot,\cdot\rangle_{X_{Z}}=\langle\cdot,\cdot\rangle_{1,Z}+\langle\cdot,\cdot\rangle$, with $\langle\cdot,\cdot\rangle_{1,Z}$ defined in (2.10). Define $\mathcal{B}=\mathcal{A}-\gamma I$, $\gamma=\beta+1>0$, with $\beta\geqq 0$ as in (2.9). Then the following linear initial value problem
    $\begin{cases}\boldsymbol{u}_{t}=\mathcal{B}\boldsymbol{u}\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}\in D(\mathcal{B})=D(\mathcal{F}_{Z})\times\mathcal{E}(\mathcal{\mathcal{G}})\end{cases}$ (2.13)
    has a unique solution $\boldsymbol{u}\in C([0,\infty):D(\mathcal{B}))\cap C^{1}([0,\infty):X_{Z})$ given by $\boldsymbol{u}(t)=\mathcal{U}(t)\boldsymbol{u}_{0}$, $t\geqq 0$, where $\{\mathcal{U}(t)\}_{t\geqq 0}$ is a $C_{0}$-semigroup of contractions on $X_{Z}$. Indeed, first we observe that $\mathcal{B}$ is dissipative on $X_{Z}$: for $\mathbf{U}=(\boldsymbol{u},\boldsymbol{v})\in D(\mathcal{B})$ with $\boldsymbol{u}=(u_{j})_{j=1}^{2}\in D(\mathcal{F}_{Z})$ and $\boldsymbol{v}=(v_{j})_{j=1}^{2}\in\mathcal{E}(\mathcal{\mathcal{G}})$ (with $u_{j},v_{j}$ real-valued without lost of generality)
    $\langle-\mathcal{B}\mathbf{U},\mathbf{U}\rangle_{X_{Z}}=(\beta+1)\Big{[}\|\boldsymbol{u}\|^{2}_{X^{1}_{Z}}+\|\boldsymbol{v}\|^{2}_{L^{2}(\mathcal{\mathcal{G}})}-\langle\boldsymbol{u},\boldsymbol{v}\rangle\Big{]}\geqq 0,$ (2.14)
    because of the Cauchy-Schwartz inequality and $\|\boldsymbol{u}\|^{2}_{L^{2}(\mathcal{\mathcal{G}})}\leqq\|\boldsymbol{u}\|^{2}_{X^{1}_{Z}}$ by (2.8). Next, we notice that from Theorem 2.3 we can choose $\lambda$ such that $\lambda+\gamma>0$ and $\lambda+\gamma\in\rho(\mathcal{A})$. Hence the range, $R(\lambda I-\mathcal{B})=R((\lambda+\gamma)I-\mathcal{A})$, of $\lambda I-\mathcal{B}$ is $X_{Z}$. Henceforth, upon application of the classical Lumer-Phillips Theorem (see, e.g., Pazy [43]) we conclude that $\mathcal{B}$ is the infinitesimal generator of a $C_{0}$-semigroup of contractions $\{\mathcal{U}(t)\}_{t\geqq 0}$ on $X_{Z}$, as claimed.
  2. (2)
    Now define $\mathcal{S}(t)=e^{\gamma t}\mathcal{U}(t)$. Then $\{\mathcal{S}(t)\}_{t\geqq 0}$ is a $C_{0}$-semigroup on $X_{Z}$ with infinitesimal generator $\mathcal{S}^{\prime}(0)=\gamma I+\mathcal{A}-\gamma I=\mathcal{A}$. Then, since the norm $\|\cdot\|_{H^{1}(\mathcal{\mathcal{G}})}$ is equivalent to the norm $\|\cdot\|_{X_{Z}^{1}}$ on $H^{1}(\mathcal{\mathcal{G}})$, we obtain that $\{\mathcal{S}(t)\}_{t\geqq 0}$ is a $C_{0}$-semigroup on $H^{1}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})$ and $\boldsymbol{w}(t)=\mathcal{S}(t)\boldsymbol{w}_{0}$ is the unique solution for the linear problem (2.11). Moreover, $\|\mathcal{S}(t)\|_{H^{1}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})}\leqq Me^{\gamma t}$, for $M>0$, $\gamma=\beta+1>0$ and $t\geqq 0$. Therefore, from Theorem 2.3, standard semigroup theory and properties of the Laplace transform, we obtain for $\theta>\beta+1$ the representation formula in (2.12). This finishes the proof.
@@
Based on Theorem 2.5 and Proposition 2.6, the local well-posedness result in $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ follows from the Banach fixed point theorem and standard arguments. It is to be noticed that the contraction mapping principle has the advantage that, since the nonlinear term $F(z)$ is smooth, then the Implicit Function Theorem implies the smoothness of the data-solution map associated to the sine-Gordon equation (see [11, 14, 15, 16] for further information). Lastly, from Proposition 2.6 we obtain that for all $t\in(0,T]$, $\boldsymbol{w}(t)\in\mathcal{E}(\mathcal{G})\times\mathcal{C}(\mathcal{Y})$. This finishes the proof. ∎
@@ 3 Linear instability criterion for the sine-Gordon model on a tadpole graph
In this section we establish a linear instability criterion of stationary solutions for the sine-Gordon model (2.2) on a tadpole graph $\mathcal{G}$. More importantly, the criterion also applies to any type of stationary solutions independently of the boundary conditions under consideration and can be therefore used to study configurations with boundary rules at the vertex of $\delta^{\prime}$-interaction type, or with other types of stationary solutions to the sine-Gordon equation.
Let us suppose that $JE$ on a domain $D(JE)\subset L^{2}(\mathcal{G})$ is the infinitesimal generator of a $C_{0}$-semigroup on $L^{2}(\mathcal{G})$ and the stationary solution $\Phi=(\phi_{1}(x),\phi_{2}(x),0,0)\in D(JE)$. Thus, every component satisfies the equation
$-c^{2}_{j}\phi^{\prime\prime}_{j}+\sin\phi_{j}=0,\quad j=1,2.$ (3.1)
Now, we suppose that $\mathbb{W}$ satisfies formally equality in (2.2) and we define
@@
with $\Psi\in D(J\mathcal{E})$. If we denote by $\sigma(J\mathcal{E})=\sigma_{\mathrm{pt}}(J\mathcal{E})\cup\sigma_{\mathrm{ess}}(J\mathcal{E})$ the spectrum of $J\mathcal{E}$ (namely, $\lambda\in\sigma_{\mathrm{pt}}(J\mathcal{E})$ if $\lambda$ is isolated and with finite multiplicity), the later discussion suggests the usefulness of the following definition.
@@
The stationary vector solution $\Phi\in D(\mathcal{E})$ is said to be spectrally stable for model sine-Gordon if the spectrum of $J\mathcal{E}$, $\sigma(J\mathcal{\mathcal{missing}}E)$, satisfies $\sigma(J\mathcal{E})\subset i\mathbb{R}.$ Otherwise, the stationary solution $\Phi\in D(\mathcal{E})$ is said to be spectrally unstable.
It is standard to show that $\sigma_{\mathrm{pt}}(J\mathcal{E})$ is symmetric with respect to both the real and imaginary axes and that $\sigma_{\mathrm{ess}}(J\mathcal{E})\subset i\mathbb{R}$, by supposing $J$ skew-symmetric and $\mathcal{E}$ self-adjoint (for instance, under Assumption $(S_{3})$ below for $\mathcal{L}$; see [28, Lemma 5.6 and Theorem 5.8]). These cases on $J$ and $\mathcal{E}$ will be considered in our theory. Hence it is equivalent to say that $\Phi\in D(J\mathcal{E})$ is spectrally stable if $\sigma_{\mathrm{pt}}(J\mathcal{E})\subset i\mathbb{R}$, and it is spectrally unstable if $\sigma_{\mathrm{pt}}(J\mathcal{E})$ contains a point $\lambda$ with $\mathrm{Re}\,\lambda>0$.
From (3.5), our eigenvalue problem to solve is now reduced to,
$J\mathcal{E}\Psi=\lambda\Psi,\quad\mathrm{Re}\,\lambda>0,\;\;\Psi\in D(\mathcal{E}).$ (3.6)
@@
Our linear instability criterion is the following.
@@
Suppose the assumptions $(S_{1})$ $(S_{3})$ hold. Then the operator $J\mathcal{E}$ has a real positive and a real negative eigenvalue.
The proof of Theorem 3.2 is based in ideas from Lopes [35], Angulo and Cavalcante [8, 9] (see also Angulo and Plaza [14] for a complete proof of this), applied to the case of a tadpole graph.
@@
A stationary vector solution $\Phi\in D(\mathcal{E})$ is said to be nonlinearly unstable in $X\equiv H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})$-norm for model sine-Gordon if there is $\epsilon>0$, such that for every $\delta>0$ there exist an initial datum $\mathbb{w}_{0}$ with $\|\Phi-\mathbb{w}_{0}\|_{X}<\delta$ and an instant $t_{0}=t_{0}(\mathbb{w}_{0})$ such that $\|\mathbb{w}(t_{0})-\Phi\|_{X}>\epsilon$, where $\mathbb{w}=\mathbb{w}(t)$ is the solution of the sine-Gordon model with $\mathbb{w}(0)=\mathbb{w}_{0}$.

### v2

@@ 2 Well-posedness for the sine-Gordon on a tadpole graph
The proof of Theorem 2.1 follows the script established in Angulo and Plaza [13] (see Section 2) in the case of the sine-Gordon model on a $\mathcal{Y}$-junction with a $\delta$-interaction. Thus, we start by describing the spectrum of the one parameter family of self-adjoint operators $(-\Delta,D_{Z})$ for $Z\in\mathbb{R}$ defined in (1.2) and (1.7) (for a related result in the context of tadpole graphs see Noja and Pelinovsky [34], Proposition A.1). In order to simplify the notation we write $\mathcal{F}_{Z}$ to denote the operator $\Delta$ acting on $L^{2}(\mathcal{G})$ with domain $D_{Z}$, for each $Z\in\mathbb{R}$.
@@
For every $Z\in\mathbb{R}$ the essential spectrum of the self-adjoint operator $\mathcal{F}_{Z}$ is purely absolutely continuous and $\sigma_{\mathrm{\tiny{ess}}}(\mathcal{F}_{Z})=\sigma_{\mathrm{\tiny{ac}}}(\mathcal{F}_{Z})=[0,\infty)$. If $Z>0$ then $\mathcal{F}_{Z}$ has exactly one negative eigenvalue, i.e., its point spectrum is $\sigma_{\mathrm{\tiny{pt}}}(\mathcal{F}_{Z})=\{-\varrho_{{}_{Z}}^{2}\}$, where $\varrho_{{}_{Z}}>0$ is the only positive root of the trascendental equation
@@
$\boldsymbol{\Phi}_{Z}=\begin{pmatrix}\cosh(\frac{\varrho_{{}_{Z}}x}{c_{1}})\\ \\ e^{-\frac{\varrho_{{}_{Z}}x}{c_{2}}}\end{pmatrix}$.
@@
First, it is to be observed that from Theorem A.6 in [6] and Proposition A.4, the one parameter family of self-adjoint extensions $(\mathcal{F}_{Z},D_{Z})$ satisfies $n(\mathcal{F}_{Z})\leqq 1$ for all $Z\in\mathbb{R}$.
@@
$\left\{\begin{aligned} -c_{1}^{2}u^{\prime\prime}&=\lambda u,\qquad x\in(-L,L)\\ u(L)&=u(-L)=0,\\ u^{\prime}(L)&-u^{\prime}(-L)=0,\end{aligned}\right$.
@@
assuming $Z>0$. Clearly, G(0)=Z<0G(0)=-Z<0 and $G(\varrho)\to\infty$ as $\varrho\to\infty$. Moreover,
@@
for all $\varrho\in(0,\infty)$ and, thus, $G$ is strictly increasing. By the intermediate value theorem there exists a unique value $\varrho_{{}_{Z}}>0$ such that $G(\varrho_{{}_{Z}})=0$.
@@
Let $Z\in\mathbb{R}$. For any $\lambda\in\mathbb{C}$ such that $\lambda^{2}\in\rho(\mathcal{F}_{Z})$ we have that $\lambda$ belongs to the resolvent set of $\mathcal{A}=JE$, with $D(\mathcal{A})=D_{Z}\times L^{2}(\mathcal{G})$; moreover, the resolvent,
@@
The proof is very similar to that of Theorem 2.2 in [13] so we omit it. ∎
It is actually possible to provide some explicit formulae for $R(-\lambda^{2}:\mathcal{F}_{Z})$ in the case where $\lambda^{2}\in\rho(\mathcal{F}_{Z})$ and $\lambda\in\mathbb{R}$. For the sake of completeness and for the reader’s convenience we now establish those.
@@
Fix $Z\in\mathbb{R}$ and consider $\lambda^{2}\in\rho(\mathcal{F}_{Z})$ with $\lambda\in\mathbb{R}$. For any $\boldsymbol{u}=(u_{1},u_{2})\in L^{2}(\mathcal{G})$ we have:
@@
    In the case when $Z\leq 0$ and for all $\lambda>0$ (without loss of generality) such that $\lambda^{2}\in\rho(\mathcal{F}_{Z})$, set $\Phi:=(\mathcal{F}_{Z}+\lambda^{2}I_{2})^{-1}\boldsymbol{u}$ to obtain
    $\Phi_{1}(x)$ $=\frac{d_{1}}{c_{1}^{2}}\cosh(\lambda x/c_{1})+\frac{1}{2c_{1}\lambda}\int_{-L}^{L}u_{1}(y)\cosh(\lambda|x-y|/c_{1})\,dy,$ $x\in(-L,L),$ (2.6)
    $\Phi_{2}(x)$ $=\frac{d_{2}}{c_{2}^{2}}e^{-\lambda x/c_{2}}+\frac{1}{2c_{2}\lambda}\int_{L}^{\infty}u_{2}(y)e^{-\lambda|x-y|/c_{2}}\,dy,$ $x\in(L,\infty).$
@@
    In the case when $Z>0$ and for all $\lambda>0$ (without loss of generality) such that $\lambda^{2}\in\rho(\mathcal{F}_{Z})$ and with $\lambda\neq\varrho_{{}_{Z}}^{2}$, set $\Psi:=(\mathcal{F}_{Z}+\lambda^{2}I_{2})^{-1}\boldsymbol{u}$ to obtain
@@
Formulae (2.6) and (2.7) result from the variation of constants formula: a solution to the homogeneous problem plus a convolution with the Green’s function associated to the operator (see formula (2.6) when $Z\leqq 0$). In the case when there exists a negative eigenvalue (that is, when $Z>0$) one also needs to project onto the associated eigenspace (formula (2.7)). The reader can directly verify that $(\mathcal{F}_{Z}+\lambda^{2}I_{2})\Phi=\boldsymbol{u}$ if $Z\leqq 0$, and that $(\mathcal{F}_{Z}+\lambda^{2}I_{2})\Psi=\boldsymbol{u}$ if $Z>0$, whenever $\lambda^{2}\in\rho(\mathcal{F}_{Z})$.
@@
$\beta:=\begin{cases}\varrho_{{}_{Z}}^{2},&\text{if }Z>0,\\ 0,&\text{if }Z\leqq 0.\end{cases}$ (2.9)
We will denote by $H^{1}_{Z}(\mathcal{G})$ the space $H^{1}(\mathcal{G})$ with the norm $\|\cdot\|_{X^{1}_{Z}}$. Moreover, for $\boldsymbol{u}=(u_{j})_{j=1}^{2}$, the following well-defined inner product in $H^{1}_{Z}(\mathcal{G})$,
$\langle\boldsymbol{u},\boldsymbol{v}\rangle_{1,Z}=\int_{-L}^{L}u^{\prime}_{1}(x)\overline{v^{\prime}_{1}(x)}dx+\int_{L}^{\infty}u^{\prime}_{2}(x)\overline{v^{\prime}_{2}(x)}dx+(\beta+1)\langle\boldsymbol{u},\boldsymbol{v}\rangle+Zu_{1}(L)\overline{v_{1}(L)},$ (2.10)
@@
The following theorem verifies that the operator $\mathcal{A}=JE$ is indeed the infinitesimal generator of a $C_{0}$-semigroup upon application of the classical Lumer-Phillips theory (see Pazy [35]).
@@
By considering the Hilbert space $X_{Z}\equiv H^{1}_{Z}(\mathcal{\mathcal{G}})\times L^{2}(\mathcal{\mathcal{G}})$ with inner product $\langle\cdot,\cdot\rangle_{X_{Z}}=\langle\cdot,\cdot\rangle_{1,Z}+\langle\cdot,\cdot\rangle$, with $\langle\cdot,\cdot\rangle_{1,Z}$ defined in (2.10), we can follow the strategy of proof of Theorem 2.5 in Angulo and Plaza [13] and from standard semigroup theory and properties of the Laplace transform (see Pazy [35]) we obtain the result.
@@
Based on Theorem 2.5 and Proposition 2.6, the local well-posedness result in $\mathcal{E}(\mathcal{G})\times L^{2}(\mathcal{G})$ follows from the Banach fixed point theorem and standard arguments. It is to be noticed that the contraction mapping principle has the advantage that, since the nonlinear term $F(\mathbf{W})$ is smooth, then the Implicit Function Theorem implies the smoothness of the data-solution map associated to the sine-Gordon equation (see [10, 13, 14, 15] for further information). Lastly, from Proposition 2.6 we obtain that for all $t\in(0,T]$, $\boldsymbol{w}(t)\in\mathcal{E}(\mathcal{G})\times\mathcal{C}(\mathcal{Y})$. This finishes the proof. ∎
@@ 3 Linear instability criterion for the sine-Gordon model on a tadpole graph
In this section we establish a linear instability criterion of stationary solutions for the sine-Gordon model (2.2) on a tadpole graph $\mathcal{G}$. It is based on the results by Angulo and Cavalcante in [7, 8] for stationary solutions of evolution models on metric graphs. By the sake of completeness we establish the main points of our criterion. More importantly, the criterion is so general that it also applies to any type of stationary solutions independently of the boundary conditions under consideration and can be therefore used to study configurations with boundary rules at the vertex of $\delta^{\prime}$-interaction type, or with other types of stationary solutions to the sine-Gordon equation.
Let us suppose that $JE$ in (2.3) on a domain $D(JE)\subset L^{2}(\mathcal{G})$ is the infinitesimal generator of a $C_{0}$-semigroup on $L^{2}(\mathcal{G})$ and we consider a priori a stationary solution $\Phi=(\phi_{1}(x),\phi_{2}(x),0,0)$ for the sine-Gordon model on a tadpole graph such that $\Phi\in D(JE)$. Thus, every component satisfies the equation
$c^{2}_{j}\phi^{\prime\prime}_{j}+\sin\phi_{j}=0,\quad j=1,2.$ (3.1)
Now, we suppose that $\mathbb{W}$ satisfies formally equality in (2.2) and we define the new variable
@@
with $\Psi\in D(J\mathcal{E})$. If we denote by $\sigma(J\mathcal{E})$ the spectrum of $J\mathcal{E}$, it is well known that we can divide this into two disjoint parts $\sigma(J\mathcal{E})=\sigma_{\mathrm{pt}}(J\mathcal{E})\cup\sigma_{\mathrm{ess}}(J\mathcal{E})$, where $\lambda\in\sigma_{\mathrm{pt}}(J\mathcal{E})$ if $\lambda$ is isolated and with finite multiplicity, and $\sigma_{\mathrm{ess}}(J\mathcal{E})$ is the essential spectrum for $J\mathcal{E}$ (see [37]). The later discussion suggests the usefulness of the following definition.
@@
The stationary vector solution $\Phi\in D(\mathcal{E})$ is said to be spectrally stable for model sine-Gordon if the spectrum of $J\mathcal{E}$, $\sigma(J\mathcal{E})$, satisfies $\sigma(J\mathcal{E})\subset i\mathbb{R}.$ Otherwise, the stationary solution $\Phi\in D(\mathcal{E})$ is said to be spectrally unstable.
It is standard to show that $\sigma_{\mathrm{pt}}(J\mathcal{E})$ is symmetric with respect to both the real and imaginary axes and that $\sigma_{\mathrm{ess}}(J\mathcal{E})\subset i\mathbb{R}$, by supposing $J$ skew-symmetric and $\mathcal{E}$ self-adjoint (for instance, under Assumption $(S_{3})$ below for $\mathcal{L}$; see [21, Lemma 5.6 and Theorem 5.8]). These cases on $J$ and $\mathcal{E}$ will be considered in our theory. Hence it is equivalent to say that $\Phi\in D(J\mathcal{E})$ is spectrally stable if $\sigma_{\mathrm{pt}}(J\mathcal{E})\subset i\mathbb{R}$, and it is spectrally unstable if $\sigma_{\mathrm{pt}}(J\mathcal{E})$ contains a point $\lambda$ with $\mathrm{Re}\,\lambda>0$.
@@
Our linear instability criterion of stationary solutions for the sine-Gordon on a tadpole graph is the following.
@@
Suppose the assumptions $(S_{1})$ - $(S_{3})$ hold. Then the operator $J\mathcal{E}$ has a real positive and a real negative eigenvalue.
The proof of Theorem 3.2 can be seen in [13] -Theorem3.2.
@@
A stationary vector solution $\Phi\in D(\mathcal{E})$ is said to be nonlinearly unstable in $X\equiv H^{1}(\mathcal{G})\times L^{2}(\mathcal{G})$-norm for model sine-Gordon (2.2) if there is $\epsilon>0$, such that for every $\delta>0$ there exist an initial datum $\mathbb{w}_{0}$ with $\|\Phi-\mathbb{w}_{0}\|_{X}<\delta$ and an instant $t_{0}=t_{0}(\mathbb{w}_{0})$ such that $\|\mathbb{W}(t_{0})-\Phi\|_{X}>\epsilon$, where $\mathbb{W}=\mathbb{W}(t)$ is the solution of the sine-Gordon model with $\mathbb{W}(0)=\mathbb{w}_{0}$.